OpenStudy (anonymous):

If n - 3 = 7 + m, what is the value of 2(n - m) + 4?

OpenStudy (anonymous):

if we subtract m from both sides we have????

OpenStudy (anonymous):

n-m=7+3 4(10)+4=44